accents: é

LaTEX

now inline: blaé test

now inline: bla étest

now inline: bla é test

real spaces xxx

tildes accent õ

horizontal 2 - is –

horizontal 3 - is —

ligatures: ffl fi fl Yp

what about vertikal bars — horizontal bar - and tildes ¡ funny carrots ¿ or other symbols:

[ ] _{^{?}=;/.,+∗()′} so far not in verbatim (tt font)

now in verbatim (tt font)
accents: \'{e}
\LaTeX
now inline: bla\'{e} test
now inline: bla \'{e}test
now inline: bla \'{e} test
real spaces~~~xxx
tildes  accent \~o
horizontal 2 -  is --
horizontal 3 -  is ---
ligatures: ffl fi fl Yp
what about vertikal bars |
horizontal bar -
and tildes ~
< funny carrots >
or other symbols: \\ { } [ ] _^? = ; / . , + * ( ) ' & % $ # " !